Grade 12Algebra

In a class test containing 10 questions, 5 marks are awarded for every correct answer, and (–2) marks are awarded for every incorrect answer and 0 for questions not attempted.
Reshma gets five correct answers and five incorrect answers, what is her score?

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

From the question,
Marks awarded for 1 correct answer = 5
Then,
Total marks awarded for 5 correct answer = 5 × 5 = 25
Marks awarded for 1 wrong answer = -2
Then,
Total marks awarded for 5 wrong answer = 5 × -2 = -10
∴Total score obtained by Reshma = 25 + (-10)
= 25–10
= 15

Thanks